# Quantum teleportation implies symmetry-protected topological order

### v2

@@ Appendix F Proof of Lemma 8
We then observe that recovery channels conditioned on other measurement outcomes also act on $\Gamma$ via Eq. F.120. Additionally, only the recovery channels $\mathcal{R}_{j,k}$ encountered prior to the measurement channel $\mathcal{M}_{j}$ (in the Heisenberg picture) are actually sensitive to the measurement outcome, and hence we convert all other “trivial” recovery channels into physical unitaries. We then define
$\Lambda_{\Gamma,j}=\sum_{k}\lambda_{j,k}~{}{\rm mod}~{}2\,,~{}~{}$ (F.121)
@@
which captures whether or not the combination of all recovery channels $\mathcal{R}_{j,k}$ conditioned on the $j$th outcome act nontrivially on the logical operator $\Gamma$ via Eq. F.120.
@@
We can then write the physical part of the time-evolved logical operator $\Gamma(t_{j})$ immediately prior to the measurement channel $\mathcal{M}_{j}$ (in the Heisenberg picture) as
$\Gamma(t_{j})=U^{\dagger}_{M}\,\Gamma\,U_{M}\,,~{}~{}$ (F.122)
@@
where $U_{M}$ is the product of all physical Clifford gates $U$ and the outcome-independent parts of all recovery-channels in $\mathcal{W}$ after the measurement channel $\mathcal{M}_{j}$ in the Schrödinger picture. Hence, it is the same unitary $U_{j}$ that appears in Eq. F.117. The Stinespring part of $\Gamma$ is
$\widetilde{\Gamma}(t_{j})=\widetilde{Z}^{\Lambda_{\Gamma,j}}_{j}\otimes\cdots\,,~{}~{}$ (F.123)
@@
where $\cdots$ captures $\widetilde{Z}$ operators attached by recovery channels conditioned on the outcomes of other measurements.
@@
Evolving $\Gamma$ (F.122) through the measurement channel $\mathcal{M}_{j}$ results in a logical operator $\Gamma$ with physical part
@@
$\Gamma(t^{\prime}_{j})$ $=\operatorname*{tr}\limits_{{\rm ss},j}\left[\,\mathcal{M}^{\dagger}_{j}~{}\Gamma(t_{j})\otimes\widetilde{Z}^{\Lambda_{\Gamma,j}}_{j}~{}\mathcal{M}_{j}\,\widetilde{P}^{(0)}_{j}\,\right]$
@@
$=\frac{1}{4}\sum\limits_{m,n=0,1}\left(\mathds{1}+(-1)^{m}\,\bar{A}_{j}\right)\Gamma(t_{j})\left(\mathds{1}+(-1)^{n}\,\bar{A}_{j}\right)$
@@
×*$~{}~{}~{}~{}~{}~{}~{}~{}\times\left\langle*\middle|0\middle|\widetilde{X}^{m}_{j}\,\widetilde{Z}^{\Lambda_{\Gamma,j}}_{j}\,\widetilde{X}^{n}_{j}\right\rangle{0}$
@@
$=\sum\limits_{n=0,1}\frac{\left(-1\right)^{n\,\Lambda_{\Gamma,j}}}{4}\left(\bar{A}_{j}+\left(-1\right)^{n}\mathds{1}\right)\left\{\bar{A}_{j},\,\Gamma(t_{j})\right\}$
$=\frac{1}{2}~{}\bar{A}^{\,\Lambda_{\Gamma,j}-1}_{j}~{}\left\{\bar{A}_{j},\,\Gamma(t_{j})\right\}\,,~{}~{}$ (F.124)
@@
$\mathds{1}$ $=\Gamma^{2}(t_{j}^{\prime})=\frac{1}{2}\mathds{1}+\frac{1}{4}\left\{\Gamma(t_{j}),\,\bar{A}_{j}\,\Gamma(t_{j})\,\bar{A}_{j}\right\}\,,~{}$ (F.125)
independent of $\Lambda_{\Gamma,j}$, leading to
@@
$=\left[\Gamma(t_{j}),\,\bar{A}_{j}\right]^{2}\,,~{}~{}$ (F.126)
@@
$B=\left[\Gamma(t_{j}),\,\bar{A}_{j}\right]=-B^{\dagger}\,,~{}~{}$ (F.127)
@@
$\left[A_{j},\,\Gamma(t_{j})\right]=0\,,~{}~{}$ (39)
@@
for all logical operators $\Gamma$ evolved to the time $t_{j}$ (F.122) immediately prior to the measurement of $A_{j}$, in the Heisenberg picture. Given Eq. 39, Eq. F.124 becomes
$\Gamma(t^{\prime}_{j})=\bar{A}^{\,\Lambda_{\Gamma,j}}_{j}~{}\Gamma(t_{j})\,,~{}~{}$ (F.128)
@@
and the same result recovers from Eq. F.124 in the case where $\Gamma(t_{j})$ has no support on site $j$. We note that the latter scenario also satisfies the commutation condition (39); hence, both Eqs. 39 and F.128 hold generally.
@@
Thus, successful teleportation requires that the measured observable $A_{j}$ commutes with all logical operators $\Gamma(t_{j})$ evolved in the Heisenberg picture to the time $t_{j}$ immediately prior to the corresponding measurement (39). Then, the combination of measuring $A_{j}$ and all recovery operations $\mathcal{R}_{j,k}$ (F.115) conditioned on the outcome attaches the involutory part $\bar{A}_{j}$ (4) of the measured observable $A_{j}$ to $\Gamma(t_{j})$ (F.122) via Eq. F.128, provided that an odd number of recovery Pauli operators $\chi_{j,k}$ anticommute with the corresponding Heisenberg-evolved logical operators $\Gamma(\tau_{j,k})$ (E.113) immediately prior to the $k$th recovery channel conditioned on the $j$th outcome—i.e., $\Lambda_{\Gamma,j}=1$ (F.121). This holds for each logical operator $\Gamma$.
@@
We now restore the Stinespring operators in the $\cdots$ term in Eq. F.123, and repeat this procedure for the remaining measurements. We note that subsequent recovery operations (in the Heisenberg picture) generally commute with $\bar{A}_{j}$, as described below Eq. F.119. The other possibility is that $\bar{A}_{j}$ is a Pauli string that anticommutes with one or more $\chi_{i,\ell}$ recovery operators (F.115); however, this remains compatible with the update in Eq. F.120, since the operator conjugated by the channel $\mathcal{R}_{i,\ell}$ is still a Pauli string (until all recovery operations have concluded, all physical unitaries are elements of $\mathbf{C}_{*}{\rm ph}$). Thus, all subsequent recovery channels $\mathcal{R}_{i,\ell}$ (in the Heisenberg picture) either act trivially on $\bar{A}_{j}$ in Eq. F.119, or $\bar{A}_{j}$ is a Pauli string that can be absorbed into $\Gamma(t_{j})$.
@@
These properties are preserved by all subsequent gates and measurements. When all recovery channels have been encountered, we allow generic (i.e., non-Clifford) physical unitaries to appear in $\mathcal{W}$, which may be absorbed into the unitary operators $U$ in Eq. F.122. Since only the conjugation by recovery channels $\mathcal{R}_{i,\ell}$ is sensitive to having a Pauli-string input, this does not pose an obstacle to evolving under any remaining measurement channels. It is straightforward to verify that the same conditions that applied to prior recovery and measurement channels also apply to all remaining dilated channels.
@@
Finally, we note that in a canonical-form protocol $\mathcal{W}=\mathcal{R}\,\mathcal{M}\,\mathcal{U}$, by Lemma 7, $\Gamma(\tau_{j})=\Gamma$ for all $j\in[1,M]$. However, the operators $\Gamma(t_{j})$ may be equal to $\Gamma$ times up to $M-j$ involutory operators, per Eq. F.128. ∎
@@ Appendix G Proof of Proposition 9
Consider the logical operators $X_{{\rm L}}$ and $Z_{{\rm L}}$, and Eq. 1 of Prop. 1. If the dilated channels attach $\bar{A}$ to $X_{{\rm L}}$ alone, then $Z_{{\rm L}}$ only grows under $\mathcal{U}$, so that $L\leq vT$ for that operator, and hence the logical qubit. The same holds if we switch $X_{{\rm L}}$ and $Z_{{\rm L}}$. If the dilated channels attach $\bar{A}$ to both $X_{{\rm L}}$ and $Z_{{\rm L}}$, then they attach $\bar{A}^{2}=\mathds{1}$ to $Y_{{\rm L}}$, which therefore obeys $L\leq vT$. Hence, a single measurement cannot be used to teleport all three logical basis operators a distance $L>vT$. For $k=M=1$, this can also be proven for any physical teleportation protocol by showing that Hermiticity of $\mathsf{S}_{\nu}$ is equivalent to $\left[\sigma^{\nu}_{i},\,\mathcal{U}^{\dagger}\,\bar{A}\,\mathcal{U}\right]=0$ for all $\nu$, which implies that at most two logical operators can exceed the Lieb-Robinson bound (28).
Next, consider a standard teleportation protocol $\mathcal{W}$ with $k\geq 1$ with $M<2k$. For $\mathcal{W}$ to succeed with fidelity one, the teleportation conditions (1) of Prop. 1 must be satisfied for all $k$ pairs of logical operators. To exceed the Lieb-Robinson bound $L\leq vT$ (28), distinct stabilizers $\mathsf{S}_{n,\nu}$ must realize for all $2k$ logical basis operators—equivalently, a different involutory operator $\bar{A}_{j}$ must be attached to all $2k$ logical basis operators. Otherwise, one can define another valid set of logical (basis) operators by taking products of logical operators and/or their associated stabilizers $\mathsf{S}_{n,\nu}$ to find one or more logical operators that satisfy $\sigma^{\nu}_{i_{n}}=\mathcal{U}^{\dagger}\,\sigma^{\nu}_{f_{n}}\,\mathcal{U}$ (1), which can only be satisfied if $L\leq d(i_{n},f_{n})\leq vT$, as with $k=1$ Friedman et al. (2022a).
@@ Appendix H Proof of Lemma 10
Again, we first consider the $k=1$ case, and generalize to $k>1$ afterward. For concreteness, consider the Heisenberg evolution of the logical operators $X_{{\rm L}}$ and $Z_{{\rm L}}$, which act on the final state $\lvert\Psi_{T}\rangle$ (7) as $X_{f}$ and $Z_{f}$, respectively. In this context, Lemma 7 guarantees that
$\mathcal{W}=\mathcal{R}_{x}\,\mathcal{R}_{z}\,\mathcal{M}_{2}\,\mathcal{M}_{1}\,\mathcal{U}\,,~{}~{}$ (H.130)
where the order of the effective recovery channels $\mathcal{R}_{x,z}$ is inconsequential, and $\mathcal{M}$ and $\mathcal{U}$ are unmodified.
@@
$\sigma^{\nu}_{\rm L}(\tau)$ $=\mathcal{R}^{\dagger}\,\sigma^{\nu}_{f}\,\mathcal{R}=\sigma^{\nu}_{f}\,\widetilde{Z}^{\lambda_{\nu,1}}_{1}\,\widetilde{Z}^{\lambda_{\nu,2}}_{2}\,,~{}~{}$ (H.131)
where $\lambda_{\nu,i}=1$ if an odd number of recovery operators $\chi_{i,k}$ based on the $i$th measurement outcome anticommute with $\sigma^{\nu}_{f}$, and is zero otherwise (for $\nu=x,y,z$). Evolving the operator $\sigma^{\nu}_{\rm L}(\tau)$ through the two measurement of $A_{2}$ (and projecting onto the default state $\lvert 0\rangle$ of the corresponding Stinespring register) leads to
$\sigma^{\nu}_{\rm L}(t_{2})$ $=\operatorname*{tr}\limits_{{\rm ss},2}\left[\,\mathcal{M}^{\dagger}_{2}\,\sigma^{\nu}_{\rm L}(\tau)\,\mathcal{M}_{2}\,\widetilde{P}^{(0)}_{2}\,\right]$
$=\sigma^{\nu}_{f}\,A^{\lambda_{\nu,2}}_{2}\,\widetilde{Z}^{\lambda_{\nu,1}}_{1}\,,~{}~{}$ (H.132)
@@
$\left[A_{2},\,\sigma^{\nu}_{\rm L}(\tau)\right]=\left[A_{2},\,\sigma^{\nu}_{f}\right]=0\,,~{}~{}$ (H.133)
@@
$\sigma^{\nu}_{\rm L}(t_{1})$ $=\operatorname*{tr}\limits_{{\rm ss},2}\left[\,\mathcal{M}^{\dagger}_{1}\,\sigma^{\nu}_{\rm L}(t_{2})\,\mathcal{M}_{1}\,\widetilde{P}^{(0)}_{1}\,\right]$
$=\sigma^{\nu}_{f}\,A^{\lambda_{\nu,2}}_{2}\,A^{\lambda_{\nu,1}}_{1}\,,~{}~{}$ (H.134)
@@
$\left[A_{1},\,\sigma^{\nu}_{\rm L}(t_{2})\right]=\left[A_{1},\,\sigma^{\nu}_{f}\,A^{\lambda_{\nu,2}}_{2}\right]=0\,,~{}~{}$ (H.135)
@@
First, $\lambda_{\nu,i}=1$ must hold for at least one logical operator labelled $\nu$, for both $i=1,2$; otherwise, one of the measurements acts trivially by Prop. 5, and the other is then useless to $\mathcal{W}$ by Prop. 9. Second, if $\lambda_{\nu,i}=1$ for both $i=1,2$ and $\nu=x,z$, e.g., then $X_{{\rm L}}(t_{1})$ and $Z_{{\rm L}}(t_{1})$ are both multiplied by $A_{2}\,A_{1}$; however, this is equivalent to a single measurement of the involutory operator $B=A_{2}\,A_{1}$, which cannot be useful to teleportation by Prop. 9. Third, if $\lambda_{\nu,i}=0$ for $i=1,2$ for any $\nu$, then that logical operator must be teleported by $\mathcal{U}$ alone, in which case $L\leq vT$, violating the conditions of the Lemma. Hence, $\mathcal{W}$ must attach one of the observables $A_{1,2}$ to $X_{{\rm L}}$ and the other to $Z_{{\rm L}}$ (or any pair of logical operators).
Suppose that $\mathcal{W}$ attaches $A_{2}$ to $Z_{{\rm L}}$ and $A_{1}$ to $X_{{\rm L}}$—i.e., $\lambda_{z,2}=\lambda_{x,1}=1$ while $\lambda_{x,2}=\lambda_{z,1}=0$. The commutation conditions imposed by Lemma 8 take the form
$\left[A_{2},\,X_{f}\right]$ $=\left[A_{1},\,X_{f}\right]=0$ (H.136a)
$\left[A_{2},\,Z_{f}\right]$ $=\left[A_{1},\,Z_{f}A_{2}\right]=0\,,~{}~{}$ (H.136b)

### v3

@@ Appendix F Proof of Lemma 8
We then observe that recovery channels conditioned on other measurement outcomes also act on $\Gamma$ via Eq. F.120. Additionally, only the recovery channels $\mathcal{R}^{\,}_{j,k}$ encountered prior to the measurement channel $\mathcal{M}^{\,}_{j}$ (in the Heisenberg picture) are actually sensitive to the measurement outcome, and hence we convert all other “trivial” recovery channels into physical unitaries. We then define
$\Lambda^{\,}_{\Gamma,j}=\sum_{k}\lambda^{\,}_{j,k}~{}{\rm mod}~{}2\,,~{}~{}$ (F.121)
@@
which captures whether or not the combination of all recovery channels $\mathcal{R}^{\,}_{j,k}$ conditioned on the $j$th outcome act nontrivially on the logical operator $\Gamma$ via Eq. F.120.
@@
We can then write the physical part of the time-evolved logical operator $\Gamma(t^{\,}_{j})$ immediately prior to the measurement channel $\mathcal{M}^{\,}_{j}$ (in the Heisenberg picture) as
$\Gamma(t^{\vphantom{\prime}}_{j})=U^{\dagger}_{M}\,\Gamma\,U^{\vphantom{\dagger}}_{M}\,,~{}~{}$ (F.122)
@@
where $U^{\,}_{M}$ is the product of all physical Clifford gates $U$ and the outcome-independent parts of all recovery-channels in $\mathcal{W}$ after the measurement channel $\mathcal{M}^{\,}_{j}$ in the Schrödinger picture. Hence, it is the same unitary $U^{\,}_{j}$ that appears in Eq. F.117. The Stinespring part of $\Gamma$ is
$\widetilde{\Gamma}(t^{\vphantom{\prime}}_{j})=\widetilde{Z}^{\Lambda^{\vphantom{\prime}}_{\Gamma,j}}_{j}\otimes\cdots\,,~{}~{}$ (F.123)
@@
where $\cdots$ captures $\widetilde{Z}^{\vphantom{*}}$ operators attached by recovery channels conditioned on the outcomes of other measurements.
@@
Evolving $\Gamma$ (F.122) through the measurement channel $\mathcal{M}^{\,}_{j}$ results in a logical operator $\Gamma$ with physical part
@@
$\Gamma(t^{\prime}_{j})$ $=\operatorname*{tr}\limits_{{\rm ss},j}\left[\,\mathcal{M}^{\dagger}_{j}~{}\Gamma(t^{\vphantom{\prime}}_{j})\otimes\widetilde{Z}^{\Lambda^{\vphantom{\prime}}_{\Gamma,j}}_{j}~{}\mathcal{M}^{\vphantom{\dagger}}_{j}\,\widetilde{P}^{(0)}_{j}\,\right]$
@@
$=\frac{1}{4}\sum\limits_{m,n=0,1}\left(\mathds{1}+(-1)^{m}\,\bar{A}^{\vphantom{\prime}}_{j}\right)\Gamma(t^{\vphantom{\prime}}_{j})\left(\mathds{1}+(-1)^{n}\,\bar{A}^{\vphantom{\prime}}_{j}\right)$
@@
×$~{}~{}~{}~{}~{}~{}~{}~{}\times\left\langle*\middle|0\middle|\widetilde{X}^{m}_{j}\,\widetilde{Z}^{\Lambda^{\vphantom{\prime}}_{\Gamma,j}}_{j}\,\widetilde{X}^{n}_{j}\right\rangle{0}$
@@
$=\sum\limits_{n=0,1}\frac{\left(-1\right)^{n\,\Lambda^{\vphantom{\prime}}_{\Gamma,j}}}{4}\left(\bar{A}^{\vphantom{\prime}}_{j}+\left(-1\right)^{n}\mathds{1}\right)\left\{\bar{A}^{\vphantom{\prime}}_{j},\,\Gamma(t^{\vphantom{\prime}}_{j})\right\}$
$=\frac{1}{2}~{}\bar{A}^{\,\Lambda^{\vphantom{\prime}}_{\Gamma,j}-1}_{j}~{}\left\{\bar{A}^{\vphantom{\prime}}_{j},\,\Gamma(t^{\vphantom{\prime}}_{j})\right\}\,,~{}~{}$ (F.124)
@@
$\mathds{1}$ $=\Gamma^{2}(t_{j}^{\prime})=\frac{1}{2}\mathds{1}+\frac{1}{4}\left\{\Gamma(t^{\vphantom{\prime}}_{j}),\,\bar{A}^{\vphantom{\prime}}_{j}\,\Gamma(t^{\vphantom{\prime}}_{j})\,\bar{A}^{\vphantom{\prime}}_{j}\right\}\,,~{}$ (F.125)
independent of $\Lambda^{\,}_{\Gamma,j}$, leading to
@@
$=\left[\Gamma(t^{\vphantom{\prime}}_{j}),\,\bar{A}^{\vphantom{\prime}}_{j}\right]^{2}\,,~{}~{}$ (F.126)
@@
$B=\left[\Gamma(t^{\vphantom{\prime}}_{j}),\,\bar{A}^{\vphantom{\prime}}_{j}\right]=-B^{\dagger}\,,~{}~{}$ (F.127)
@@
$\left[A^{\vphantom{\prime}}_{j},\,\Gamma(t^{\vphantom{\prime}}_{j})\right]=0\,,~{}~{}$ (39)
@@
for all logical operators $\Gamma$ evolved to the time $t^{\,}_{j}$ (F.122) immediately prior to the measurement of $A^{\,}_{j}$, in the Heisenberg picture. Given Eq. 39, Eq. F.124 becomes
$\Gamma(t^{\prime}_{j})=\bar{A}^{\,\Lambda^{\vphantom{\prime}}_{\Gamma,j}}_{j}~{}\Gamma(t^{\vphantom{\prime}}_{j})\,,~{}~{}$ (F.128)
@@
and the same result recovers from Eq. F.124 in the case where $\Gamma(t^{\,}_{j})$ has no support on site $j$. We note that the latter scenario also satisfies the commutation condition (39); hence, both Eqs. 39 and F.128 hold generally.
@@
Thus, successful teleportation requires that the measured observable $A^{\,}_{j}$ commutes with all logical operators $\Gamma(t^{\,}_{j})$ evolved in the Heisenberg picture to the time $t^{\,}_{j}$ immediately prior to the corresponding measurement (39). Then, the combination of measuring $A^{\,}_{j}$ and all recovery operations $\mathcal{R}^{\,}_{j,k}$ (F.115) conditioned on the outcome attaches the involutory part $\bar{A}^{\,}_{j}$ (33) of the measured observable $A^{\,}_{j}$ to $\Gamma(t^{\,}_{j})$ (F.122) via Eq. F.128, provided that an odd number of recovery Pauli operators $\chi^{\,}_{j,k}$ anticommute with the corresponding Heisenberg-evolved logical operators $\Gamma(\tau^{\,}_{j,k})$ (E.113) immediately prior to the $k$th recovery channel conditioned on the $j$th outcome—i.e., $\Lambda^{\,}_{\Gamma,j}=1$ (F.121). This holds for each logical operator $\Gamma$.
@@
We now restore the Stinespring operators in the $\cdots$ term in Eq. F.123, and repeat this procedure for the remaining measurements. We note that subsequent recovery operations (in the Heisenberg picture) generally commute with $\bar{A}^{\,}_{j}$, as described below Eq. F.119. The other possibility is that $\bar{A}^{\,}_{j}$ is a Pauli string that anticommutes with one or more $\chi^{\,}_{i,\ell}$ recovery operators (F.115); however, this remains compatible with the update in Eq. F.120, since the operator conjugated by the channel $\mathcal{R}^{\,}_{i,\ell}$ is still a Pauli string (until all recovery operations have concluded, all physical unitaries are elements of $\mathbf{C}^{\vphantom{*}}_{*}{\rm ph}$). Thus, all subsequent recovery channels $\mathcal{R}^{\,}_{i,\ell}$ (in the Heisenberg picture) either act trivially on $\bar{A}^{\,}_{j}$ in Eq. F.119, or $\bar{A}^{\,}_{j}$ is a Pauli string that can be absorbed into $\Gamma(t^{\,}_{j})$.
@@
These properties are preserved by all subsequent gates and measurements. When all recovery channels have been encountered, we allow generic (i.e., non-Clifford) physical unitaries to appear in $\mathcal{W}$, which may be absorbed into the unitary operators $U$ in Eq. F.122. Since only the conjugation by recovery channels $\mathcal{R}^{\,}_{i,\ell}$ is sensitive to having a Pauli-string input, this does not pose an obstacle to evolving under any remaining measurement channels. It is straightforward to verify that the same conditions that applied to prior recovery and measurement channels also apply to all remaining dilated channels.
@@
Finally, we note that in a canonical-form protocol $\mathcal{W}=\mathcal{R}\,\mathcal{M}\,\mathcal{U}$, by Lemma 7, $\Gamma(\tau^{\,}_{j})=\Gamma$ for all $j\in[1,M]$. However, the operators $\Gamma(t^{\,}_{j})$ may be equal to $\Gamma$ times up to $M-j$ involutory operators, per Eq. F.128. ∎
@@ Appendix G Proof of Proposition 9
Consider the logical operators $X^{\vphantom{*}}_{{\rm L}}$ and $Z^{\vphantom{*}}_{{\rm L}}$, and Eq. 8 of Prop. 1. If the dilated channels attach $\bar{A}$ to $X^{\vphantom{*}}_{{\rm L}}$ alone, then $Z^{\vphantom{*}}_{{\rm L}}$ only grows under $\mathcal{U}$, so that $L\leq vT$ for that operator, and hence the logical qubit. The same holds if we switch $X^{\vphantom{*}}_{{\rm L}}$ and $Z^{\vphantom{*}}_{{\rm L}}$. If the dilated channels attach $\bar{A}$ to both $X^{\vphantom{*}}_{{\rm L}}$ and $Z^{\vphantom{*}}_{{\rm L}}$, then they attach $\bar{A}^{2}=\mathds{1}$ to $Y^{\vphantom{*}}_{{\rm L}}$, which therefore obeys $L\leq vT$. Hence, a single measurement cannot be used to teleport all three logical basis operators a distance $L>vT$. For $k=M=1$, this can also be proven for any physical teleportation protocol by showing that Hermiticity of $\mathsf{S}^{\,}_{\nu}$ is equivalent to $\left[\sigma^{\nu}_{i},\,\mathcal{U}^{\dagger}\,\bar{A}\,\mathcal{U}\right]=0$ for all $\nu$, which implies that at most two logical operators can exceed the Lieb-Robinson bound (28).
Next, consider a standard teleportation protocol $\mathcal{W}$ with $k\geq 1$ with $M<2k$. For $\mathcal{W}$ to succeed with fidelity one, the teleportation conditions (8) of Prop. 1 must be satisfied for all $k$ pairs of logical operators. To exceed the Lieb-Robinson bound $L\leq vT$ (28), distinct stabilizers $\mathsf{S}^{\,}_{n,\nu}$ must realize for all $2k$ logical basis operators—equivalently, a different involutory operator $\bar{A}^{\,}_{j}$ must be attached to all $2k$ logical basis operators. Otherwise, one can define another valid set of logical (basis) operators by taking products of logical operators and/or their associated stabilizers $\mathsf{S}^{\,}_{n,\nu}$ to find one or more logical operators that satisfy $\sigma^{\nu}_{i_{n}}=\mathcal{U}^{\dagger}\,\sigma^{\nu}_{f_{n}}\,\mathcal{U}$ (8), which can only be satisfied if $L\leq d(i_{n},f_{n})\leq vT$, as with $k=1$ [42].
@@ Appendix H Proof of Lemma 10
Again, we first consider the $k=1$ case, and generalize to $k>1$ afterward. For concreteness, consider the Heisenberg evolution of the logical operators $X^{\vphantom{*}}_{{\rm L}}$ and $Z^{\vphantom{*}}_{{\rm L}}$, which act on the final state $\lvert\Psi^{\,}_{T}\rangle$ (7) as $X^{\vphantom{*}}_{f}$ and $Z^{\vphantom{*}}_{f}$, respectively. In this context, Lemma 7 guarantees that
$\mathcal{W}=\mathcal{R}^{\vphantom{\dagger}}_{x}\,\mathcal{R}^{\vphantom{\dagger}}_{z}\,\mathcal{M}^{\vphantom{\dagger}}_{2}\,\mathcal{M}^{\vphantom{\dagger}}_{1}\,\mathcal{U}\,,~{}~{}$ (H.130)
where the order of the effective recovery channels $\mathcal{R}^{\,}_{x,z}$ is inconsequential, and $\mathcal{M}$ and $\mathcal{U}$ are unmodified.
@@
$\sigma^{\nu}_{\rm L}(\tau)$ $=\mathcal{R}^{\dagger}\,\sigma^{\nu}_{f}\,\mathcal{R}=\sigma^{\nu}_{f}\,\widetilde{Z}^{\lambda^{\vphantom{\prime}}_{\nu,1}}_{1}\,\widetilde{Z}^{\lambda^{\vphantom{\prime}}_{\nu,2}}_{2}\,,~{}~{}$ (H.131)
where $\lambda^{\vphantom{\prime}}_{\nu,i}=1$ if an odd number of recovery operators $\chi^{\,}_{i,k}$ based on the $i$th measurement outcome anticommute with $\sigma^{\nu}_{f}$, and is zero otherwise (for $\nu=x,y,z$). Evolving the operator $\sigma^{\nu}_{\rm L}(\tau)$ through the two measurement of $A^{\,}_{2}$ (and projecting onto the default state $\lvert 0\rangle$ of the corresponding Stinespring register) leads to
$\sigma^{\nu}_{\rm L}(t^{\,}_{2})$ $=\operatorname*{tr}\limits_{{\rm ss},2}\left[\,\mathcal{M}^{\dagger}_{2}\,\sigma^{\nu}_{\rm L}(\tau)\,\mathcal{M}^{\vphantom{\dagger}}_{2}\,\widetilde{P}^{(0)}_{2}\,\right]$
$=\sigma^{\nu}_{f}\,A^{\lambda^{\vphantom{\prime}}_{\nu,2}}_{2}\,\widetilde{Z}^{\lambda^{\vphantom{\prime}}_{\nu,1}}_{1}\,,~{}~{}$ (H.132)
@@
$\left[A^{\vphantom{\prime}}_{2},\,\sigma^{\nu}_{\rm L}(\tau)\right]=\left[A^{\vphantom{\prime}}_{2},\,\sigma^{\nu}_{f}\right]=0\,,~{}~{}$ (H.133)
@@
$\sigma^{\nu}_{\rm L}(t^{\,}_{1})$ $=\operatorname*{tr}\limits_{{\rm ss},2}\left[\,\mathcal{M}^{\dagger}_{1}\,\sigma^{\nu}_{\rm L}(t^{\,}_{2})\,\mathcal{M}^{\vphantom{\dagger}}_{1}\,\widetilde{P}^{(0)}_{1}\,\right]$
$=\sigma^{\nu}_{f}\,A^{\lambda^{\vphantom{\prime}}_{\nu,2}}_{2}\,A^{\lambda^{\vphantom{\prime}}_{\nu,1}}_{1}\,,~{}~{}$ (H.134)
@@
$\left[A^{\vphantom{\prime}}_{1},\,\sigma^{\nu}_{\rm L}(t^{\,}_{2})\right]=\left[A^{\vphantom{\prime}}_{1},\,\sigma^{\nu}_{f}\,A^{\lambda^{\vphantom{\prime}}_{\nu,2}}_{2}\right]=0\,,~{}~{}$ (H.135)
@@
First, $\lambda^{\,}_{\nu,i}=1$ must hold for at least one logical operator labelled $\nu$, for both $i=1,2$; otherwise, one of the measurements acts trivially by Prop. 5, and the other is then useless to $\mathcal{W}$ by Prop. 9. Second, if $\lambda^{\,}_{\nu,i}=1$ for both $i=1,2$ and $\nu=x,z$, e.g., then $X^{\vphantom{*}}_{{\rm L}}(t^{\,}_{1})$ and $Z^{\vphantom{*}}_{{\rm L}}(t^{\,}_{1})$ are both multiplied by $A^{\,}_{2}\,A^{\,}_{1}$; however, this is equivalent to a single measurement of the involutory operator $B=A^{\,}_{2}\,A^{\,}_{1}$, which cannot be useful to teleportation by Prop. 9. Third, if $\lambda^{\,}_{\nu,i}=0$ for $i=1,2$ for any $\nu$, then that logical operator must be teleported by $\mathcal{U}$ alone, in which case $L\leq vT$, violating the conditions of the Lemma. Hence, $\mathcal{W}$ must attach one of the observables $A^{\,}_{1,2}$ to $X^{\vphantom{*}}_{{\rm L}}$ and the other to $Z^{\vphantom{*}}_{{\rm L}}$ (or any pair of logical operators).
Suppose that $\mathcal{W}$ attaches $A^{\,}_{2}$ to $Z^{\vphantom{*}}_{{\rm L}}$ and $A^{\,}_{1}$ to $X^{\vphantom{*}}_{{\rm L}}$—i.e., $\lambda^{\,}_{z,2}=\lambda^{\,}_{x,1}=1$ while $\lambda^{\,}_{x,2}=\lambda^{\,}_{z,1}=0$. The commutation conditions imposed by Lemma 8 take the form
$\left[A^{\vphantom{\prime}}_{2},\,X^{\vphantom{*}}_{f}\right]$ $=\left[A^{\vphantom{\prime}}_{1},\,X^{\vphantom{*}}_{f}\right]=0$ (H.136a)
$\left[A^{\vphantom{\prime}}_{2},\,Z^{\vphantom{*}}_{f}\right]$ $=\left[A^{\vphantom{\prime}}_{1},\,Z^{\vphantom{*}}_{f}A^{\vphantom{\prime}}_{2}\right]=0\,,~{}~{}$ (H.136b)